# Detecting Dataset Bias in Medical AI: A Generalized and Modality-Agnostic Auditing Framework

### v1

@@ Abstract
Data-driven AI is establishing itself at the center of evidence-based medicine. However, reports of shortcomings and unexpected behavior are growing due to AI’s reliance on association-based learning. A major reason for this behavior: latent bias in machine learning datasets can be amplified during training and/or hidden during testing. We present a data modality-agnostic auditing framework for generating targeted hypotheses about sources of bias which we refer to as Generalized Attribute Utility and Detectability-Induced bias Testing (G-AUDIT) for datasets. Our method examines the relationship between task-level annotations and data properties including protected attributes (e.g., race, age, sex) and environment and acquisition characteristics (e.g., clinical site, imaging protocols). G-AUDIT automatically quantifies the extent to which the observed data attributes may enable shortcut learning, or in the case of testing data, hide predictions made based on spurious associations. We demonstrate the broad applicability and value of our method by analyzing large-scale medical datasets for three distinct modalities and learning tasks: skin lesion classification in images, stigmatizing language classification in Electronic Health Records (EHR), and mortality prediction for ICU tabular data. In each setting, G-AUDIT successfully identifies subtle biases commonly overlooked by traditional qualitative methods that focus primarily on social and ethical objectives, underscoring its practical value in exposing dataset-level risks and supporting the downstream development of reliable AI systems. Our method paves the way for achieving deeper understanding of machine learning datasets throughout the AI development life-cycle from initial prototyping all the way to regulation, and creates opportunities to reduce model bias, enabling safer and more trustworthy AI systems.
@@ 1 Introduction
The use of AI in healthcare necessitates methods and algorithms that ensure safe, unbiased, and fair outcomes. Recent shifts towards data-driven methods, where performance depends heavily on increasingly large datasets, has created pressure to rapidly scale data collection and curation efforts. Collecting large datasets that meet data quality and diversity requirements for fair and safe algorithm development remains an ongoing challenge.
Datasets collected for machine learning (ML) applications often need to (1) accurately represent populations of interest [41], (2) ensure fair treatment and minimize bias wherever possible [40, 47, 26, 21, 16, 1], and (3) represent the types of conditions/quality expected in real-world settings while avoiding the inclusion of learning shortcuts [32, 49, 48, 50, 34]. These requirements are further impacted by privacy, safety, cost, and other systematic constraints that hinder or prevent the ability to ensure all desired dataset properties are met [24, 46, 12].
Oftentimes, data collection constraints and requirements may be at odds or impractical to achieve. For instance, some populations may not be well-represented by fixed-sample datasets (even when those datasets are considered large-scale) [2, 14, 35]. Similarly, balancing datasets for parity across protected subgroups may create unintentional biases relating to the representation of particular clinical sites, imaging protocols, or other unprotected groups[14, 37]. Even the notions of fairness and bias are nuanced, disconcertingly suggesting that data collection cannot possibly satisfy all definitions [4]. In short, despite careful efforts to construct unbiased datasets, dataset bias may be inevitable and may manifest in forms beyond traditional ethical or social considerations [24].
@@
Without effective methods to audit and correct dataset bias, data-driven AI methods are at considerable risk of learning and amplifying implicit biases. Recent research shows that this risk manifests through fairness/bias in AI predictions [15, 41, 40, 16, 21], shortcut learning [48, 49, 34, 22, 9], and lack of generalization or robustness [36] all of which may result in harmful performance disparities of the model within and across patient populations. While algorithmic auditing methods [27, 1] provide the means to detect, quantify, and characterize limitations of trained machine learning models, they are applied post-hoc and force mitigation strategies to be reactive rather than proactive. The best opportunity to address data-driven AI bias risks starts with exposing data-level bias prior to model training or evaluation. Therefore, data audits, the focus of this work, are of utmost importance for identifying and reducing model bias as early as possible in the training and evaluation phases.
@@
Figure 1: Given a dataset consisting of input with associated labels and metadata, we conduct our Generalized Attribute Utility and Detectability-Induced bias Testing (G-AUDIT) for data procedure to generate quantitative hypotheses on the relative risk of attributes in the form of a Detectability vs Utility scatter plot (right).
@@
Despite the importance of identifying bias at the dataset-level, techniques for performing dataset audits remain largely absent [24]. On one hand, some methods may examine disparities in medical dataset metadata [21, 1, 35] but they do not explicitly link these to the primary data with which machine learning models are trained. In contrast, other methods show that patient attributes may be predictable from sensor-level measurements [15, 17] but may not fully identify whether the attribute-level features represent direct shortcuts for achieving accurate prediction. Some methods more directly examine the relation between data attributes and task model accuracy [41, 40, 22] yet these do not enable assessment of the dataset-level risks independent of task-level modeling assumptions. Lastly, while most methods have focused on addressing bias with respect to conventional protected attributes (such as race, age, sex), some methods have begun to explore biases that may relate to other aspects of the data collection pipeline such as image acquisition or clinician-level marking [37, 38]. In short, current dataset auditing methods do not holistically address both the risk of relevant attributes leaking information about the task and the risk of whether information about those attributes can be directly exploited during model training. Automated dataset auditing techniques are required to address these primary risks and generate hypotheses to guide downstream model development, auditing, and mitigation.
@@
We present a novel technique that represents a significant advance in enabling independent, dataset-level auditing and which we refer to as Generalized Attribute Utility and Detectability-Induced bias Testing (G-AUDIT) for datasets. Our method presents the first unified approach to shortcut auditing by considering the interplay between attribute-level composition, sensor-level measurements, and task labels. Our method is not limited to only protected patient-level attributes and allows for datasets to be audited with respect to any variables of relevance to the data collection process, patient population, and targeted machine learning task.
Our auditing procedure (Figure 1) utilizes the tools of information theory and causal inference to identify the presence and strength of bias between attributes ($A$) and task labels ($Y$) (utility) as well as the ability to infer the values of those attributes directly from the data ($X$) itself (detectability). In particular, we first quantify utility as the strength of the association between $A\leftrightarrow Y$. However, high utility is not sufficient for demonstrating the presence of a shortcut because we also need to estimate how easily information about $A$ can be inferred from $X$ to solve the task. We therefore next measure detectability as the association $A\leftrightarrow X$ by predicting $\hat{A}$ from $X$ while controlling for the influence of $Y$. High-utility and -detectability are indicative of potential shortcuts and would be flagged by our auditing method as targets for downstream model audits or review by clinicians. In this work, we consider the issue of dataset bias across a diverse range of health applications and domains including tabular, text, and image data and identify sources of bias commonly overlooked or previously unknown.
@@ 2 Results
To demonstrate its broad effectiveness, our G-AUDIT procedure is applied to datasets from three different modalities, namely, image, text, and tabular data. For each dataset, we estimate the utility and detectability (see Sec. 4) of each attribute relative to the underlying learning task. We do this to generate hypotheses that identify and rank potential shortcut attributes according to the risk that downstream models may exploit them. We also include results from an optional calibration procedure in which an approximate upper bound on the drop in performance-related metrics can be calculated for each potential shortcut using a synthetic attribute in precisely controlled conditions (see Section 4.4). The bound provides a means of estimating a form of worst case downstream model performance risks for specific attributes in more familiar metrics.
@@ 2.1 Skin Lesion Classification
We first consider shortcut risks in vision-based tasks. The high dimensionality of image data creates many opportunities for shortcuts to exist without directly impacting the machine learning task. For instance, hospital-specific tokens placed in the chest X-ray field of view (e.g., [50]) may impact only a small number of pixels in the image but may constitute a shortcut when associated primarily with a specific disease condition in the training dataset. In cases like these, DNNs trained on the data may exploit salient features like tokens/watermarks (or other statistical regularities not task-related) to achieve low training error. We focus here on skin lesion classification where the construction of large-scale datasets may not be able to adequately balance across patient characteristics, clinical sites, and dermascopic imaging sensors/settings and where such features of the dataset and collection process may manifest as shortcuts.
The ISIC 2019 skin lesion dataset [43, 6, 7] was analyzed for bias with respect to the included metadata and patient attributes. The dataset consists of 25,331 training images and attributes age, race, sex, anatomical location, and skin color on the Fitzpatrick scale. Image metadata includes height, width, and year of collection. While the original task labels included seven diagnostic categories, we reduce the classification task to malignant vs. benign conditions (e.g., [51, 29, 31]). The data auditing procedure (see Sec. 4) was applied to the entire training dataset relative to the binary malignancy classification task and excluding any images with missing attribute values.
The main auditing results are found in Figure 2 where the height, width, and year attributes exhibit the highest combination of utility and detectability, indicating these attributes are more likely sources of bias within the dataset. While these particular attributes may seem unrelated to the task labels, this image metadata can act as a proxy for the camera type and/or clinical site. Furthermore, while all images are resized and cropped to a fixed resolution prior to running the auditing procedure, the high detectability scores indicate that some of this information is still retained in the images themselves, either directly or via proxy.
Figure 2: (left) Attribute Detectability vs. Utility for the ISIC 2019 dataset. (right) Approximate worst-case performance risk for synthetic attribute with varying utility. Vertical lines indicate observed values of utility and indicate a worse-case drop in AUC of $\approx 0.2$.
Table 1: Comparison between Utility and Detectability measures as well as macro-averaged F1 scores from the detection model and conducting the SPLIT procedure [15, 17] on trained task models.
@@ 2.3 Mortality Prediction with Intensive Care Unit Tabular Data
Lastly, we consider the dataset auditing problem in the context of tabular data. While not as feature rich or high dimensional as images and text, tabular data presents its own unique challenges from an auditing perspective. While the features learned by task models in the image and text domains often lack interpretability (e.g., DNN embeddings), tabular data provides a direct mapping between attributes and their measured values. In these cases, attributes used as inputs to a task model can still act as shortcuts when they exhibit undesirable associations with the task labels that could arise due to sample bias or incorrect usage of the attribute itself.
For instance, if clinical sites differ in their test-ordering protocols, associations between disease conditions and clinical site may be reflected in the test orders/results provided to machine learning task models [42]. Furthermore, the gap in task performance between DNNs and more traditional ML models (e.g., SVMs, ensemble methods, etc.) is much smaller in the tabular case and this provides an opportunity to measure both detectability and task model performance over a wider class of machine learning models.
@@
Figure 5: We order each attribute from least to greatest recoverability in a downstream FT-Transformer task model as measured by SPLIT [17, 15]. By doing this, we see that attributes that are more recoverable tend to have higher detectability values, though there is variance from model to model.
@@
Figure 6: Following our calibration procedure, we can explore the worst case impact a completely detectable shortcut has on downstream models. Dotted vertical lines correspond to the utility values of real attributes in the dataset. Green lines correspond to test sets where the shortcut’s relationship to the label holds true; red correspond to the worst-case counterfactual set.
@@ 3 Discussion
G-AUDIT is the first known method for fully quantitative measurement of dataset bias relative to both clinical and imaging factors. Prior to this method, the closest points of comparison consider only protected attributes (such as race, age, or sex) and overlook other aspects of the dataset which may be a stronger source of bias (such as the collection site or model of imaging device). In fact, our analysis finds that non-patient attributes are often of greatest concern which is consistent with previous studies that focused on specific shortcut learning scenarios such as the use of digital watermarks in x-ray images [9] or presence of chest drains for detecting pneumothorax [35]. While these cases were found through manual inspection by researchers, our auditing procedure would automatically and more efficiently enable detection of these possible sources of bias prior to model training. Since our audits are implicitly tied to measurable and interpretable attributes of the data, clinicians and AI developers are better equipped to interpret the results of the audit and identify relevant courses of action (e.g., bias mitigation or model auditing strategies).
We show the generality of G-AUDIT by applying it across multiple machine learning tasks and data modalities. G-AUDIT is shown to be effective in both causal and anti-causal scenarios where the relationship between the labels and data may be reversed. While G-AUDIT requires domain knowledge to estimate the directionality, our method gracefully handles each condition given the a priori assumption. Future work will consider augmenting G-AUDIT to include detection of label-data directionality.
@@
In this work, we provide a generalized, quantitative technique for generating hypotheses about dataset bias to inform downstream model training and auditing. As adoption of data-driven machine learning methods for safety- and cost-critical medical applications continues to grow, we must have principled quantitative methods for analyzing the underlying training and evaluation data. G-AUDIT provides the first such approach and results demonstrate that commonly overlooked dataset attributes may induce dataset bias and ultimately lead to AI failures and disparities in diagnosis and treatment of underlying patient conditions. Our method provides a positive step towards identifying and mitigating these risks.
@@ 4.1 Utility and Detectability
For our purposes, datasets consist of samples ($X$), associated metadata or attributes ($A$), and task labels ($Y$). We establish the existence (not direction) of a relationship between $A\leftrightarrow Y$ through a measure we call utility. The utility of an attribute refers to our ability to infer the value of the task label $Y$ simply by observing $A$. In the extreme case, if $A$ is perfectly correlated with $Y$, then machine learning models simply need to detect $A$ in order to correctly solve the task.
However, a large value for an attribute’s utility is not sufficient to consider it as a shortcut. For this, we require to know the attribute’s detectability. Since machine learning models will typically only take $X$ as input, detectability measures the extent to which the value of $A$ can be inferred from $X$.
Attributes with high-utility and high-detectability represent the greatest risk for biasing downstream task models. However, while we use utility and detectability as a proxy for risk, features that are causally relevant to the task and have high utility and detectability are useful, rather than representing possible shortcuts. We must rely on domain expertise to determine whether high-risk attributes are reasonable features or unanticipated dataset flaws.

### v2

@@ Abstract
Artificial Intelligence (AI) is now firmly at the center of evidence-based medicine. Despite many success stories that edge the path of AI’s rise in healthcare, there are comparably many reports of significant shortcomings and unexpected behavior of AI in deployment. A major reason for these limitations is AI’s reliance on association-based learning, where non-representative machine learning datasets can amplify latent bias during training and/or hide it during testing. To unlock new tools capable of foreseeing and preventing such AI bias issues, we present G-AUDIT. Generalized Attribute Utility and Detectability-Induced bias Testing (G-AUDIT) for datasets is a modality-agnostic auditing framework that allows for generating targeted hypotheses about sources of bias in training or testing data. Our method examines the relationship between task-level annotations (commonly referred to as “labels”) and data properties including patient attributes (e.g., age, sex) and environment and acquisition characteristics (e.g., clinical site, imaging protocols). G-AUDIT automatically quantifies the extent to which the observed data attributes pose a risk for shortcut learning, or in the case of testing data, might hide predictions made based on spurious associations. We demonstrate the broad applicability and value of our method by analyzing large-scale medical datasets for three distinct modalities and machine learning tasks: skin lesion classification in images, stigmatizing language classification in Electronic Health Records (EHR), and mortality prediction for ICU tabular data. In each setting, G-AUDIT successfully identifies subtle biases commonly overlooked by traditional qualitative methods, underscoring its practical value in exposing dataset-level risks and supporting the downstream development of reliable AI systems.
@@ 1 Introduction
The use of artificial intelligence (AI) and machine learning (ML) in healthcare has created numerous opportunities to improve quality, safety, efficiency, efficacy, affordability, and accessibility of healthcare through applications including disease detection [31], forecasting [2, 24], opportunistic screening [43], and quantitative medicine [11]. However, AI/ML models have repeatedly exhibited deterioration in performance and implicit biases when moving from the development phase to testing and deployment phases [54, 37, 12, 26, 41, 15, 40]. Due to the data-driven nature of AI/ML methods, the brittleness and bias of these models is inherently tied to the data on which they are trained and evaluated prior to deployment. Bias in medical datasets may originate from a number of sources [29, 17] and the growing size of these datasets [4, 28, 27] increases the risk that latent biases may go undetected during data collection and then subsequently exploited by downstream AI/ML models.
@@
The influence of dataset biases may manifest in a number of ways including shortcut learning [52, 53, 38, 25, 9, 41], bias and fairness issues in AI predictions [16, 46, 45, 18, 23], and lack of generalization or robustness [40]. All of these issues are problematic, as they may result in harmful and unpredictable performance disparities of the model within and across patient populations in deployment. The best opportunity to address data-driven AI/ML bias risks starts with exposing dataset-level risks prior to model training or evaluation. Therefore, dataset audits, the focus of this work, are of critical importance for identifying and quantifying model bias as early as possible in the training and evaluation phases.
@@
Figure 1: Given a dataset consisting of input with associated labels and metadata, we perform our Generalized Attribute Utility and Detectability-Induced bias Testing (G-AUDIT) to generate quantitative hypotheses on the relative risk of attributes in the form of a Detectability vs Utility scatter plot (right).
@@
Despite the importance of quantifying bias risk at the dataset-level, techniques for performing dataset audits remain largely absent [29]. On the one hand, some methods may examine disparities in medical dataset metadata [23, 1, 39] but they do not explicitly link these to the primary data with which AI/ML models are trained. In contrast, other methods show that patient attributes may be predictable from sensor-level measurements [16, 19] but may not fully identify whether the attribute-level features represent direct shortcuts for achieving accurate prediction. Some methods more directly examine the relation between data attributes and task model accuracy [46, 45, 25, 48] yet these do not enable assessment of the dataset-level risks independent of task-level modeling assumptions. Lastly, while most methods have focused on addressing bias with respect to conventional patient attributes (such as age, sex), some methods have begun to explore biases that may relate to other aspects of the data collection pipeline such as image acquisition or clinician-level markings [41, 42]. In short, current dataset auditing methods lack practical utility because they do not holistically address both the risk of relevant attributes leaking information about the task and the risk of whether information about those attributes can be directly exploited during model training. Furthermore, existing auditing methods have been tailored to specific modalities or clinical tasks leading to a paucity of modality-agnostic methods for quantifying dataset-level bias risks. Thus, new automated dataset auditing techniques are required to identify these primary risks and generate hypotheses to guide downstream model development, auditing, and mitigation.
@@
To address the need for quantitative dataset audits, we present a novel technique that represents a significant advance in enabling independent, dataset-level auditing and which we refer to as Generalized Attribute Utility and Detectability-Induced bias Testing (G-AUDIT) for datasets. Our method presents the first unified approach to shortcut auditing by considering the interplay between attribute-level composition, sensor-level measurements, and task labels. G-AUDIT is generalizable and not tied to a particular data modality or ML task. Our procedure (Figure 1) utilizes the tools of information theory and causal inference to identify the presence and strength of association between attributes and task labels (utility) as well as the ability to infer the values of those attributes directly from the data itself (detectability). We then rank attributes according to their measured utility and detectability, thus providing AI/ML developers and clinicians alike a quantitative yet interpretable means of assessing dataset-level bias risk. Our method is not limited to only patient-level attributes and allows for datasets to be audited with respect to any variables of relevance to the data collection process, patient population, and machine learning task.
We apply our G-AUDIT method to a diverse set of clinical tasks covering image, text, and tabular data modalities. In each dataset, our audits uncover shortcut risks (confirmed by domain experts) linked to non-patient attributes representing aspects of the data collection pipeline not typically considered as candidates for bias. We also provide an optional calibration procedure which allows for estimating a form of worst case performance degradation given the observed values of utility for attributes in the dataset. For the potential shortcuts flagged by our auditing method, we find that the estimated worst case degradation is substantial (often $>0.2$ Area Under the Curve difference) indicating that models exploiting these shortcuts may stand to exhibit significant drops in performance in deployment. By applying G-AUDIT across clinical tasks and data modalities, our results underscore and address the need for new methods to quantitatively assess dataset bias risk across all aspects of the data collection process.
@@ 2 Results
To demonstrate its broad effectiveness, our G-AUDIT procedure is applied to datasets from three different modalities, namely, image, text, and tabular data. For each dataset, we estimate the utility and detectability (see Sec. 4) of each attribute relative to the underlying learning task. We do this to generate hypotheses that identify and rank potential shortcut attributes according to the risk that downstream models may exploit them. We also include results from an optional calibration procedure in which an approximate upper bound on the drop in performance-related metrics can be calculated for each potential shortcut using a synthetic attribute in precisely controlled conditions (see Section 4.4). The bound provides a means of estimating a form of worst-case downstream model performance risks for specific attributes in more familiar metrics.
@@ 2.1 Skin Lesion Classification
We first consider shortcut risks in vision-based tasks. The high dimensionality of image data creates many opportunities for shortcuts to exist without directly impacting the machine learning task. For instance, hospital-specific tokens placed in the chest X-ray field of view (e.g., [55]) may impact only a small number of pixels in the image but may constitute a shortcut when associated primarily with a specific disease condition in the training dataset. In cases like these, deep neural networks (DNNs) trained on the data may exploit salient features like tokens/watermarks (or other statistical regularities not task-related) to achieve low training error. We focus here on skin lesion classification where the construction of large-scale datasets may not be able to adequately balance across patient characteristics, clinical sites, and dermascopic imaging sensors/settings and where such features of the dataset and collection process may manifest as shortcuts.
The ISIC 2019 skin lesion dataset [49, 6, 7] was analyzed for bias with respect to the included metadata and patient attributes. The dataset consists of 25,331 training images and attributes age, race, sex, anatomical location, and skin color on the Fitzpatrick scale. Image metadata includes height, width, and year of collection. While the original task labels included seven diagnostic categories, we reduce the classification task to malignant vs. benign conditions (e.g., [56, 33, 35]). The data auditing procedure (see Sec. 4) was applied to the entire training dataset relative to the binary malignancy classification task. Images with any missing attribute values were excluded from the analysis.
The main auditing results are found in Figure 2 and Table 1 where the height, width, and year attributes exhibit the highest combination of utility and detectability, indicating these attributes are more likely sources of bias within the dataset. While these particular attributes may seem unrelated to the task labels, this image metadata can act as a proxy for the camera type and/or clinical site. Furthermore, while all images are resized and cropped to a fixed resolution prior to running the auditing procedure, the high detectability scores indicate that some of this information is still retained in the images themselves, either directly or via proxy.
Figure 2: (left) Attribute Detectability vs. Utility for the ISIC 2019 dataset. Confidence intervals calculated via basic bootstrap [13] (right) Approximate worst-case performance risk for a synthetic attribute with varying utility (see Sec 4.4 for more detail). Confidence intervals are derived from a t-distribution based on model performance across three folds; vertical lines indicate observed values of utility and indicate a worse-case drop in AUC of $\approx 0.2$.
Table 1: Comparison between Utility and Detectability measures as well as macro-averaged F1 scores from the detection model and conducting the SPLIT procedure [16, 19] on trained task models (see Sec. 4.5.4).
@@ 2.3 Mortality Prediction with Intensive Care Unit Tabular Data
Lastly, we consider the dataset auditing problem in the context of tabular data. While not as feature rich or high dimensional as images and text, tabular data presents its own unique challenges from an auditing perspective. While the features learned by task models in the image and text domains often lack interpretability (e.g., DNN embeddings), tabular data provides a direct mapping between attributes and their measured values. In these cases, attributes used as inputs to a task model can still act as shortcuts when they exhibit undesirable associations with the task labels that could arise due to sample bias or incorrect usage of the attribute itself. For instance, if clinical sites differ in their test-ordering protocols, associations between disease conditions and clinical site may be reflected in the test orders/results provided to machine learning task models [47]. Furthermore, the gap in task performance between DNNs and more traditional ML models (e.g., SVMs, ensemble methods, etc.) is much smaller in the tabular case and this provides an opportunity to measure both detectability and task model performance over a wider class of machine learning models.
@@
Figure 5: We order each attribute from least to greatest recoverability in a downstream FT-Transformer task model as measured by SPLIT [19, 16]. By doing this, we see that attributes that are more recoverable tend to have higher detectability values, though there is variance model to model. Additionally, we find that a large proportion of the most detectable potential shortcuts consist of missing values.
@@
Figure 6: Following our calibration procedure, we can explore the worst-case impact a completely detectable shortcut has on downstream models. Dotted vertical lines correspond to the utility values of real attributes in the dataset. Blue lines correspond to test sets where the shortcut’s relationship to the label holds true; red correspond to the worst-case counterfactual set.
@@ 3 Discussion
G-AUDIT is the first known method for fully quantitative measurement of dataset bias relative to both clinical and imaging factors. Prior to this method, the closest points of comparison consider only patient attributes (such as age or sex) and overlook other aspects of the dataset which may be a stronger source of bias (such as the collection site or model of imaging device). In fact, our analysis finds that non-patient attributes are often of greatest concern which is consistent with previous studies that focused on specific shortcut learning scenarios such as the use of digital watermarks in x-ray images [9] or presence of chest drains for detecting pneumothorax [39]. While these cases were found through manual inspection by researchers, our auditing procedure would automatically and more efficiently enable detection of these possible sources of bias prior to model training. Since our audits are implicitly tied to measurable and interpretable attributes of the data, clinicians and AI developers are better equipped to interpret the results of the audit and identify relevant courses of action (e.g., bias mitigation or model auditing strategies).
We show the generality of G-AUDIT by applying it across multiple machine learning tasks and data modalities. G-AUDIT is shown to be effective in both causal and anti-causal scenarios where the relationship between the labels and data may be reversed. While we can run G-AUDIT easily for both scenarios, additional knowledge or a priori assumptions about the directionality can be leveraged to improve the overall efficiency. Future work will consider augmenting G-AUDIT to also include automated detection of label-data directionality.
@@
In this work, we provide a generalized, quantitative technique for generating hypotheses related to the non-representativeness of a dataset to inform downstream model training and auditing. As adoption of data-driven machine learning methods for safety- and cost-critical medical applications continues to grow, principled quantitative methods for analyzing the underlying training and evaluation data are important tools for highlighting potential limitations of the data. G-AUDIT provides the first such approach and results demonstrate that commonly overlooked dataset attributes may induce biased AI outputs that negatively impact the diagnosis and treatment of underlying patient conditions. Our method provides a positive step towards identifying and mitigating these risks.
@@ 4.1 Utility and Detectability
For our purposes, datasets consist of samples ($X$), associated metadata/attributes ($A$), and task labels ($Y$). We establish the existence (not direction) of a relationship between $A\leftrightarrow Y$ through a measure we call utility. The utility of an attribute refers to our ability to infer the value of the task label $Y$ simply by observing $A$. In the extreme case, if $A$ is perfectly correlated with $Y$, then machine learning models simply need to detect $A$ in order to correctly solve the task.
However, a large value for an attribute’s utility is not sufficient to consider it a shortcut risk. For this, we require to know the attribute’s detectability. Since machine learning models will typically only take $X$ (but not $A$) as input, detectability measures the extent to which the value of $A$ can be inferred from $X$.
Attributes with high-utility and high-detectability represent the greatest risk for biasing downstream task models. However, while we use utility and detectability as a proxy for risk, features that are causally relevant to the task and have high-utility and -detectability are useful, rather than representing possible shortcuts. Therefore, domain expertise is still needed to determine whether high-risk attributes are reasonable features or unanticipated dataset flaws.